OpenStudy (anonymous):

two urns contain white balls and yellow balls. the first urn contains 5 white balls and 8 yellow balls and the second urn contains 9 white balls and 2 yellow balls. a ball drawn at random from each urn. what is the probability that both are white

OpenStudy (anonymous):

since these are independent events you multiply the probabilities together. \[\frac{5}{13}\times \frac{9}{11}\]

OpenStudy (anonymous):

First probability = 5/13 Second probability = 9/11 Both = 5/13*9/11 = 31.5%

OpenStudy (anonymous):

vvSusan keeps track of the number of tickets sold for each play presented at The Community Theater. Within how many standard deviations from the mean do all the values fall? 135, 71, 69, 80, 158, 152, 161, 96, 122, 118, 87, 85 A. 3 B. 4 C. 5 D. 2

OpenStudy (anonymous):

@tkhunny help me please

OpenStudy (tkhunny):

Why are you hijacking this tread? 1) Calculate the standard deviation. 2) Note the min and max. 3) Calculate the answer to the question.

OpenStudy (anonymous):

for the question about the urn it's 1/9